OpenStudy (anonymous):

how do you solve 2=\sqrt{3b-2}-\sqrt{10-b}\]

OpenStudy (nikvist):

\[2=\sqrt{3b-2}-\sqrt{10-b}\quad,\quad b\in[2/3,10]\] \[4=3b-2+10-b-2\sqrt{(3b-2)(10-b)}\] \[2\sqrt{(3b-2)(10-b)}=2b+4\] \[\sqrt{-3b^2+32b-20}=b+2\] \[-3b^2+32b-20=b^2+4b+4\] \[0=4b^2-28b+24=4(b-1)(b-6)\] \[b=6\]